Red goes 50 mph, must go 40 miles and starts right away. Yellow goes 40 mph, must go 30 miles and starts right away. Blue goes 75 mph, must go 50 miles, and starts 15 minutes late. Green goes 60 mph, must go 25 miles, and starts 15 minutes late. White goes 30 mph, must go 10 miles, and starts 30 minutes late. What is the order each color gets to the town?
Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

As cute as this problem is, there is not enough information, since you didn't tell us anything about the town! Nor where the colors were heading nor where they are now.
If all the colors were heading toward town.
rate * time=distance
red
50*(t)=200
t = 4
yellow
40*(t)=30
t=0.75
blue
75*(t-15)=50
t=15.6667
Green
60*(t-15)=25
t=15.4167
white
30*(t-30)=10
t=30.3333
The lowest t's arrive first

Answer by stanbon(75887)

red goes 50 mph, must go 40 miles and starts right away. Yellow goes 40 mph, must go 30 miles and starts right away. Blue goes 75 mph, must go 50 miles, and starts 15 minutes late. Green goes 60 mph, must go 25 miles, and starts 15 minutes late. White goes 30 mph, must go 10 miles, and starts 30 minutes late. What is the order each color gets to the town?
------------
For each color use time = distance/rate
The color with the smallest time will get to town first.
===========================================================
Cheers,
Stan H.